OpenStudy (anonymous):

how do input a fraction in this site (through equation) like 2*5 --- 1*5 where the three dashes represent the dash line that separates the numerator and denominator

OpenStudy (anonymous):

When you click the equation button, type \frac{2 times 5}{1 times 3}

OpenStudy (anonymous):

through equation editor

OpenStudy (anonymous):

\[\frac{2 \times 5}{1\times 5}\]